Doing some past papers and I can't find any examples of this and dont remember covering it in class, could anyone help me so I have an example going forwards Simplify the following Boolean function given in terms of Monterrey inbo pos format with a minimum number of literada Fla,b,c
Answer #1

The minimal POS of given finction is

F = (b+c')(b'+c)

Solved using k map as in image

Flaib, c) = =TIM (1,2,5,6) given solving using K-maps. f be ol 10 F = (b+7)(+c) 3 This is the mininal POS Solution. (b+c) (bt